# A Unifying Framework for Parallelizing Sequential Models with Linear Dynamical Systems

### v1

@@ D.1 Discussion of parallel chord methods
Let $\mathcal{A}$ be a fixed-point operator with fixed-point $\mathbf{x}^{*}$. Let $C(\mathcal{A},\mathbf{x}^{*})$ be the set of all sequences generated by $\mathcal{A}$ which converge to $\mathbf{x}^{*}$. Then the $R_{1}$-factors of $\mathcal{A}$ at $\mathbf{x}^{*}$ are given by
$R_{1}(\mathcal{A},\mathbf{x}^{*}):=\sup\left\{\limsup_{i\to\infty}\|\mathbf{x}^{(i)}-\mathbf{x}^{*}\|^{1/i}\,\middle|\,\{\mathbf{x}^{(i)}\}_{i\geq 0}\in C(\mathcal{A},\mathbf{x}^{*})\right\}.$ (19)
Intuitively, $R_{1}(\mathcal{A},\mathbf{x}^{*})$ gives the rate of linear convergence of a fixed-point operator $\mathcal{A}$ to its fixed-point $\mathbf{x}^{*}$. Theorem 10.1.4 of Ortega & Rheinboldt (1970) implies that if $\mathcal{A}$ is a parallel-chord method, then $R_{1}(\mathcal{A},\mathbf{x}^{*})=\sigma(\mathcal{A},F,\mathbf{x}^{*})$. Therefore, if $\sigma>0$, then $\sigma$ is the rate of $R$-linear convergence of $\mathcal{A}$ to $\mathbf{x}^{*}$, while if $\sigma=0$, we say that $\mathcal{A}$ converges $R$-superlinearly. However, it is important to note that these definitions are asymptotic in nature.
@@
Let $F(\mathbf{x})=0$ be a nonlinear equation of the form eq.˜3 with solution $\mathbf{x}^{*}$. Let $\mathcal{A}$ be a parallel-chord method with fixed-point $\mathbf{x}^{*}$. Then
$\sigma\left(\mathcal{A},F,\mathbf{x}^{*}\right)=0.$
@@
As shown in eq.˜14 and eq.˜17, both $\nicefrac{{\partial F}}{{\partial\mathbf{x}}}$ and $T$ are lower-triangular matrices with all $D\times D$ identity matrices on their main block-diagonal. In particular, $T^{-1}$ is also a lower-triangular matrix with all $D\times D$ identity matrices on its main block-diagonal. Consequently, the product $T^{-1}\nicefrac{{\partial F}}{{\partial\mathbf{x}}}$ is also a lower-triangular matrix with all $D\times D$ identity matrices on its main block-diagonal. As a result, $I-T^{-1}\nicefrac{{\partial F}}{{\partial\mathbf{x}}}$ is a lower-triangular matrix with all zeros on its main block-diagonal, and so has all its eigenvalues equal to 0. Consequently, its spectral radius is equal to zero. ∎
It may seem counterintuitive that even Jacobi iterations technically enjoy $R$-superlinear convergence in the context of parallelizing Markov processes. However, this seemingly strange result stems from the asymptotic nature of Definition˜3 of $R$-convergence, and the fact that Proposition 1 of Gonzalez et al. (2024) guarantees that all fixed-point iterations of the form given by eq.˜4 will converge to $\mathbf{x}^{*}$ in a finite number of iterations ($T$, to be exact). Therefore, for any LDS fixed-point scheme, we always have $\lim_{i\to\infty}\|\mathbf{x}^{(i)}-\mathbf{x}^{*}\|=0$.
Because this asymptotic notion of $R$-convergence is not suitable in our setting, we instead use empirical case studies in Section˜3 to show that the closeness of $A_{t}$ to $\nicefrac{{\partial f_{t}}}{{\partial x}}$ impacts the number of iterations needed for $\mathcal{A}$ to converge. This empirical approach also highlights how the increased computational cost of higher-order fixed-point methods affects wall-clock time on GPUs.
D.2 Other parallel fixed-point techniques
@@
For example, Gonzalez et al. (2024) introduces the scale-ELK algorithm which for some $k\in[0,1]$ sets the transition matrix $A$ to be
$A_{t}=(1-k)\dfrac{\partial f_{t}}{\partial x}.$
Scale-ELK can also be applied to the diagonal approximation for quasi-Newton methods. However, scale-ELK introduces an additional hyperparameter $k$, which is ideally chosen to keep the eigenvalues of $A_{t}$ below $1$ in magnitude to assure that the LDS is stable.
Therefore, we propose clip-ELK, which is a hyperparameter free approach to achieve the same goal of a stable LDS. Clip-ELK applies to the diagonal approximation only, and simply clips each element of $A_{t}$ (which in this setting is a diagonal matrix) to be between [1,1][-1,1]. Clip-ELK is immediately also a part of the framework set out in Table 1 of fixed-point methods that parallelize the evaluation of sequences via iterative application of LDS; moreover, it too must also converge globally by Proposition 1 of Gonzalez et al. (2024). An interesting and important direction for future work includes both developing and interpreting more fixed-point methods in the context of this framework that we propose.

### v2

@@ F.1 Discussion of parallel chord methods
Let $\mathcal{A}$ be a fixed-point operator with fixed-point $\mathbf{x}^{\star}$. Let $C(\mathcal{A},\mathbf{x}^{\star})$ be the set of all sequences generated by $\mathcal{A}$ which converge to $\mathbf{x}^{\star}$. Then the $R_{1}$-factors of $\mathcal{A}$ at $\mathbf{x}^{\star}$ are given by
$R_{1}(\mathcal{A},\mathbf{x}^{\star}):=\sup\left\{\limsup_{i\to\infty}\|\mathbf{x}^{(i)}-\mathbf{x}^{\star}\|^{1/i}\,\middle|\,\{\mathbf{x}^{(i)}\}_{i\geq 0}\in C(\mathcal{A},\mathbf{x}^{\star})\right\}.$ (29)
Intuitively, $R_{1}(\mathcal{A},\mathbf{x}^{\star})$ gives the rate of linear convergence of a fixed-point operator $\mathcal{A}$ to its fixed-point $\mathbf{x}^{\star}$. Theorem 10.1.4 of Ortega and Rheinboldt (1970) implies that if $\mathcal{A}$ is a one-step stationary method with update given by $\mathbf{U}(\cdot)$, then $R_{1}(\mathcal{A},\mathbf{x}^{\star})=\sigma(\mathbf{U},\mathbf{F},\mathbf{x}^{\star})$. Therefore, if $\sigma>0$, then $\sigma$ is the rate of $R$-linear convergence of $\mathcal{A}$ to $\mathbf{x}^{\star}$, while if $\sigma=0$, we say that $\mathcal{A}$ converges $R$-superlinearly. However, it is important to note that these definitions are asymptotic in nature.
@@
Let $\mathbf{F}(\mathbf{x})=\mathbf{0}$ be a nonlinear equation of the form eq.˜3 with solution $\mathbf{x}^{\star}$. Let $\mathcal{A}$ be a parallel-chord method with fixed-point $\mathbf{x}^{\star}$. Then
$\sigma\left(\mathbf{U},\mathbf{F},\mathbf{x}^{\star}\right)=0.$
@@
Both $\nicefrac{{\partial\mathbf{F}}}{{\partial\mathbf{x}}}(\mathbf{x}^{\star})$ and $\widetilde{\mathbf{J}}(\mathbf{x}^{\star})$ are lower-triangular matrices with all $D\times D$ identity matrices on their main block-diagonal. In particular, $\widetilde{\mathbf{J}}^{-1}$ is also a lower-triangular matrix with all $D\times D$ identity matrices on its main block-diagonal. Consequently, the product $\widetilde{\mathbf{J}}^{-1}\dfrac{\partial\mathbf{F}}{\partial\mathbf{x}}$ is also a lower-triangular matrix with all $D\times D$ identity matrices on its main block-diagonal. As a result, $\mathbf{I}-\widetilde{\mathbf{J}}^{-1}\dfrac{\partial\mathbf{F}}{\partial\mathbf{x}}$ is a lower-triangular matrix with all zeros on its main block-diagonal, and so has all its eigenvalues equal to 0. Consequently, its spectral radius is equal to zero. ∎
It may seem counterintuitive that even Jacobi iterations technically enjoy $R$-superlinear convergence in the context of parallelizing Markov processes. However, this seemingly strange result stems from the asymptotic nature of Definition˜3 of $R$-convergence, and the fact that Proposition 1 of Gonzalez et al. (2024) guarantees that all fixed-point iterations of the form given by eq.˜4 will converge to $\mathbf{x}^{\star}$ in a finite number of iterations ($T$, to be exact). Therefore, for any LDS fixed-point scheme, we always have $\lim_{i\to\infty}\|\mathbf{x}^{(i)}-\mathbf{x}^{\star}\|=0$.
However, in both Proposition 4 of Lu et al. (2025) and Proposition˜3 of this paper, we effectively get around this difficulty by considering the spectral norm instead of the spectral radius. The spectral norm always bounds the spectral radius, and so by focusing on spectral radius, Ortega and Rheinboldt (1970) could get tighter bounds (faster rates of convergence). However, in our setting the spectral radius cannot distinguish between any of the fixed-point methods, and so we instead use the looser bound provided by the spectral norm, which can distinguish between the different fixed-point methods. Note that the core entities are effectively the same, as $\gamma$ defined in eq.˜15 is equal to $\|\nicefrac{{\partial\mathbf{U}}}{{\partial\mathbf{x}}}(\mathbf{x}^{\star})\|_{2}$.
Finally, again, because all of our fixed-point methods converge in at most $T$ iterations, asymptotic notions of linear convergence are not suitable to fully capture the behavior of these fixed point methods. For this reason, we use empirical case studies in Section˜4 to show the efficacy of the intuition, inspired by Proposition˜3, that the closeness of $\tilde{A}_{t}$ to $\nicefrac{{\partial f_{t}}}{{\partial x}}$ impacts the number of iterations needed for $\mathcal{A}$ to converge. This empirical approach also highlights how the increased computational cost of higher-order fixed-point methods affects wall-clock time on GPUs.
F.2 Trust-region fixed-point methods
@@
For example, Gonzalez et al. (2024) introduces the scale-ELK algorithm which for some $k\in[0,1]$ sets the transition matrix $\tilde{A}_{t}$ to be
$\tilde{A}_{t}=(1-k)\dfrac{\partial f_{t+1}}{\partial x}.$
Scale-ELK can also be applied to the diagonal approximation for quasi-Newton methods. However, scale-ELK introduces an additional hyperparameter $k$.
Therefore, we propose clip-ELK, which is a hyperparameter-free approach to achieve the same goal of a stable LDS. Clip-ELK applies to the diagonal approximation only, and simply clips each element of $\tilde{A}_{t}$ (which in this setting is a diagonal matrix) to be between [1,1][-1,1]. Clip-ELK is immediately also a part of the framework set out in Table 1 of fixed-point methods that parallelize the evaluation of sequences via iterative application of LDS; moreover, it too must also converge globally by Proposition 1 of Gonzalez et al. (2024). An interesting and important direction for future work includes both developing and interpreting more fixed-point methods in the context of this framework that we propose.